OpenStudy (anonymous):

Simplify the following: 2x-squared - 2/ 6x + 6 * 6x-squared + 18x/ x-squared + 2x - 3 ?

OpenStudy (anonymous):

\[2 x^2 - \frac{1}{3}x + 36 x^2+ \frac{18}{x} + 2 x - 3=38 x^2+\frac{5 x}{3}+\frac{18}{x}-3 \]

OpenStudy (anonymous):

Uh... Im pose to end up still with a ?/?